# Construction of Solutions with Extraordinary Gradient Amplification and Localization for Schrödinger Equations

### v1

@@ 2. Preliminaries
We adopt the following notation:
$\nabla u=\nabla_{\mathbf{x}}u=\left(\partial_{x_{1}}u,\partial_{x_{2}}u,\dots,\partial_{x_{d}}u\right)=\left(u_{x_{1}},u_{x_{2}},\dots,u_{x_{d}}\right),\partial_{t}u=u_{t}=u^{\prime},\partial_{t}^{2}u=u_{tt}=u^{\prime\prime},$
and similarly for higher-order derivatives. The symbols $\Re u$ and $\Im u$ denote the real and imaginary parts of $u$, respectively, and $\mathrm{i}$ denotes the imaginary unit $\sqrt{-1}$. The floor function is written as $\lfloor x\rfloor$.
For an open set $U\subseteq\mathbb{R}^{n}$, an integer $k\geq 0$, and $0<\gamma\leq 1$, the Hölder space $C^{k,\gamma}(\bar{U})$ consists of all functions $u\in C^{k}(\bar{U})$ for which the norm
$\|u\|_{C^{k,\gamma}(\bar{U})}=\sum_{|\alpha|\leq k}\sup_{\bar{U}}|D^{\alpha}u|+\sum_{|\alpha|=k}[D^{\alpha}u]_{C^{\gamma}(\bar{U})}$
is finite, where $v]_{C^{\gamma}(\bar{U})}$ denotes the $\gamma$-Hölder seminorm:
$v]_{C^{\gamma}(\bar{U})}=\sup_{\begin{subarray}{c}\mathbf{x},\mathbf{y}\in U\\ \mathbf{x}\neq\mathbf{y}\end{subarray}}\frac{|v(\mathbf{x})-v(\mathbf{y})|}{|\mathbf{x}-\mathbf{y}|^{\gamma}}.$
Let $1\leq p\leq\infty$ and let $X$ be a Banach space. The Bochner–Sobolev space $W^{1,p}(0,T;X)$ consists of all functions $u\in L^{p}(0,T;X)$ such that the first weak derivative $u^{\prime}$ exists and belongs to $L^{p}(0,T;X)$. The norm on $W^{1,p}(0,T;X)$ is defined as
$\|u\|_{W^{1,p}(0,T;X)}=\begin{cases}\left(\int_{0}^{T}\|u(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime}(t)\|_{X}^{p}dt\right)^{\frac{1}{p}}&\text{if }1\leq p<\infty,\\ \operatorname{ess\,sup}_{0\leq t\leq T}\left(\|u(t)\|_{X}+\|u^{\prime}(t)\|_{X}\right)&\text{if }p=\infty.\end{cases}$
Similarly, the Bochner Sobolev space $W^{2,p}(0,T;X)$ consists of all functions $u\in L^{p}(0,T;X)$ such that both the first and second weak derivatives $u^{\prime}$ and $u^{\prime\prime}$ exist and belong to $L^{p}(0,T;X)$. The norm on $W^{2,p}(0,T;X)$ is given by
$\|u\|_{W^{2,p}(0,T;X)}=\begin{cases}\left(\int_{0}^{T}\|u(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime}(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime\prime}(t)\|_{X}^{p}dt\right)^{\frac{1}{p}}&\text{if }1\leq p<\infty,\\ \operatorname{ess\,sup}_{0\leq t\leq T}\left(\|u(t)\|_{X}+\|u^{\prime}(t)\|_{X}+\|u^{\prime\prime}(t)\|_{X}\right)&\text{if }p=\infty.\end{cases}$
Equipped with these norms, both $W^{1,p}(0,T;X)$ and $W^{2,p}(0,T;X)$ are Banach spaces.
@@
(see [23]) Let $\mathcal{O}$ be a bounded domain of class $C^{\alpha,1}$, $\alpha\in\mathbb{N}\cup\{0\}$, in $\mathbb{R}^{d}$. Denote by $\mathbb{H}$ the space of all Herglotz functions of the form (2.4). Define
$\mathbb{H}(\mathcal{O}):=\left\{\left.u\right|_{\mathcal{O}}:u\in\mathbb{H}\right\}$
@@
$\mathfrak{H}(\mathcal{O}):=\left\{u\in C^{\infty}(\mathcal{O}):\Delta u+\omega^{2}u=0\text{ in }\mathcal{O}\right\}.$
Then $\mathbb{H}(\mathcal{O})$ is dense in $\mathfrak{H}(\mathcal{O})\cap H^{\alpha+1}(\mathcal{O})$ with respect to the $H^{\alpha+1}(\mathcal{O})$-norm.
@@ 3. A general framework for specifying inputs
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}+L_{1}\mathcal{U}=F_{1}&\text{in }\Omega\times(0,T],\\ \mathcal{U}(\mathbf{x},0)=0&\text{in }\Omega,\\ \mathcal{U}(\mathbf{x},t)=0&\text{on }\partial\Omega\times[0,T],\end{cases}$ (4.1)
@@
$F_{1}(\mathbf{x},t)=\nabla\cdot\left((\mathbf{I}-\mathbf{A}_{1})\nabla u_{0}\right).$ (4.2)
Recalling that $\mathbf{A}_{1}-\mathbf{I}$ has compact support in $D$, we obtain
@@
Furthermore, combining (3.17), (3.18), (4.2), and the assumption $\mathbf{A}_{1}\in W^{2,\infty}(\Omega)$, we obtain the following estimate for $F_{1}$:
@@
Here, the spatial derivatives of $u_{0}$ appearing in $F_{1}$ are of order at most two, whereas the space–time Sobolev norm estimate requires third-order spatial derivatives of $H_{g}$ due to the presence of spatial gradients in the $H^{1}(\Omega)$ norm. Hence, the $H^{3}$ norm of $H_{g}$ suffices to control all terms in the estimate. The constant $C$ depends on $d$, $T$, $D$, and $\|\mathbf{A}_{1}\|_{W^{2,\infty}(\Omega)}$.
@@
    where the constant $C_{3}$ depends on where the constant $C_{3}$ depends on $d$, $T$, $\theta$, $D$, $\Omega^{\prime}$, $\Omega$, $\|\mathbf{A}_{1}\|_{W^{2,\infty}(\Omega)}$$\|\mathbf{A}_{1}\|_{C^{1,1}(\overline{\Omega})}$.
@@
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}_{1}+L_{1}\mathcal{U}_{1}=F_{1}^{\prime},&\text{in }\Omega\times(0,T],\\[5.69054pt] \mathcal{U}_{1}(x,0)=-\mathrm{i}F_{1}(x,0),&\text{in }\Omega,\\[5.69054pt] \mathcal{U}_{1}(x,t)=0,&\text{on }\partial\Omega\times[0,T].\end{cases}$ (4.9)
@@
$L_{1}\mathcal{U}(t)=F_{1}(t)-\mathrm{i}\,\partial_{t}\mathcal{U}(t)\in H^{1}(\Omega)$
in the weak sense for almost every $t\in[0,T]$. Since $\mathbf{A}_{1}\in C^{1,1}(\overline{\Omega})$, interior elliptic regularity (Theorem 8.10 in [9]) implies that, for any $\Omega^{\prime}\Subset\Omega\setminus\overline{D}$,
@@
    $\hat{L}_{2}\mathscr{U}:=\nabla\cdot(\mathbf{A}_{2}\nabla\mathscr{U})+\hat{c}\,\mathscr{U}.$
@@
    $F_{2}(\mathbf{x},t)=\nabla\cdot\big(\mathbf{I}-\mathbf{A}_{2}\big)\nabla u_{0}-c\,u_{0}-\sum_{k=2}^{l_{0}}\alpha_{k}u_{0}^{k}.$ (5.3)
@@
    where the constant $C$ depends only on $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$.
@@
Properties 1 follows immediately from the compact support in $D$ of $\mathbf{I}-\mathbf{A}_{2}(\cdot,t)$, $c(\cdot,t)$, $\alpha_{k}(\cdot,t)$ for $t\in[0,T]$, $k=2,\dots,l_{0}$ combined with the expressions defining $\hat{c}$, $\hat{N}(\mathscr{U})$ and $F_{2}$.
@@
There exists a time $T>0$ ((given below by (5.21))), depending on $\varepsilon$, $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, such that problem (5.1) admits a unique solution $\mathscr{U}$ satisfying:
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{U}+\hat{L}_{2}\mathscr{U}=F_{2}-\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.6)
@@
where constant $C_{4}(T)$ depends continuously on $T$, $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, but not on $\varepsilon$. Moreover, $C_{4}(T)$ is monotonically increasing in $T$ with $C_{4}(0)<\infty$.
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+\hat{L}_{2}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+[\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]\mathscr{U}=\partial^{\beta}_{\mathrm{x}}F_{2}-\partial^{\beta}_{\mathrm{x}}\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\partial^{\beta}_{\mathrm{x}}\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d},\end{aligned}\right$. (5.8)
where $\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]=\partial_{\mathbf{x}}^{\beta}\hat{L}_{2}-\hat{L}_{2}\partial_{\mathbf{x}}^{\beta}$ denotes the commutator, which is a differential operator of order at most 2 in spatial derivatives. Multiply the first line in (5.8) with $\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}$, use integration by parts one obtains,
$\mathrm{i}\int_{\mathbb{R}^{d}}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}-\int_{\mathbb{R}^{d}}a^{ij}_{2}(t)\partial_{i}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial_{j}\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ $+\int_{\mathbb{R}^{d}}[\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ (5.9)
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{W}+\hat{L}_{2}\mathscr{W}=-\big(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2})\big)&&\text{in }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{W}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.16)
@@
where $\partial_{\mathbf{x}}^{\gamma},\hat{L}_{2}]=\partial_{\mathbf{x}}^{\gamma}\hat{L}_{2}-\hat{L}_{2}\partial_{\mathbf{x}}^{\gamma}$ denotes the commutator, which is a differential operator of order 0 in spatial derivatives. By using (5.2), the fact that $\mathscr{V}_{1},\mathscr{V}_{2}\in X_{\varepsilon}$ implies $\|\mathscr{V}_{i}\|_{W^{1,\infty}(\mathbb{R}^{d})}\leqslant C\sqrt{\varepsilon}$ for $i=1,2$ and Lemma 5.1, we calculate the term $\partial^{\gamma}_{\mathrm{x}}(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2}))||^{2}_{L^{2}(\mathbb{R}^{d})}$ as follows,

### v2

@@ 2. Preliminaries
(see [22]) Let $\mathcal{O}$ be a bounded domain of class $C^{\alpha,1}$, $\alpha\in\mathbb{N}\cup\{0\}$, in $\mathbb{R}^{d}$. Denote by $\mathbb{T}$ the space of all Herglotz functions of the form (2.4). Define
$\mathbb{T}(\mathcal{O}):=\left\{\left.V\right|_{\mathcal{O}}:V\in\mathbb{H}\right\}$
@@
$\mathfrak{H}(\mathcal{O}):=\left\{u\in C^{\infty}(\mathcal{O}):\Delta V+\kappa^{2}V=0\text{ in }\mathcal{O}\right\},$
where $\kappa\in\mathbb{R}_{+}$. Then $\mathbb{T}(\mathcal{O})$ is dense in $\mathbb{T}(\mathcal{O})\cap H^{\alpha+1}(\mathcal{O})$ with respect to the $H^{\alpha+1}(\mathcal{O})$-norm.
@@ 3. A general framework for specifying inputs
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}+\mathcal{L}_{1}\mathcal{U}=F_{1}&\text{in }\Omega\times(0,T],\\ \mathcal{U}(\mathbf{x},0)=0&\text{in }\Omega,\\ \mathcal{U}(\mathbf{x},t)=0&\text{on }\partial\Omega\times[0,T],\end{cases}$ (4.1)
@@
$F_{1}(\mathbf{x},t)=\nabla\cdot\left((\mathbf{I}-\mathscr{A}_{1})\nabla u_{0}\right).$ (4.2)
Recalling that $\mathscr{A}_{1}-\mathbf{I}$ has compact support in $D$, we obtain
@@
Furthermore, combining (3.17), (3.18), (4.2), and the assumption $\mathscr{A}_{1}\in W^{2,\infty}(\Omega)$, we obtain the following estimate for $F_{1}$:
@@
Here, the spatial derivatives of $u_{0}$ appearing in $F_{1}$ are of order at most two, whereas the space–time Sobolev norm estimate requires third-order spatial derivatives of $H_{g}$ due to the presence of spatial gradients in the $H^{1}(\Omega)$ norm. Hence, the $H^{3}$ norm of $H_{g}$ suffices to control all terms in the estimate. The constant $C$ depends on $d$, $T$, $D$, and $\|\mathscr{A}_{1}\|_{W^{2,\infty}(\Omega)}$.
@@
    where the constant $C_{3}$ depends on where the constant $C_{3}$ depends on $d$, $T$, $\theta$, $D$, $\Omega^{\prime}$, $\Omega$, $\|\mathscr{A}_{1}\|_{W^{2,\infty}(\Omega)}$$\|\mathscr{A}_{1}\|_{C^{1,1}(\overline{\Omega})}$.
@@
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}_{1}+\mathcal{L}_{1}\mathcal{U}_{1}=F_{1}^{\prime},&\text{in }\Omega\times(0,T],\\[5.69054pt] \mathcal{U}_{1}(x,0)=-\mathrm{i}F_{1}(x,0),&\text{in }\Omega,\\[5.69054pt] \mathcal{U}_{1}(x,t)=0,&\text{on }\partial\Omega\times[0,T].\end{cases}$ (4.9)
@@
$\mathcal{L}_{1}\mathcal{U}(t)=F_{1}(t)-\mathrm{i}\,\partial_{t}\mathcal{U}(t)\in H^{1}(\Omega)$
in the weak sense for almost every $t\in[0,T]$. Since $\mathscr{A}_{1}\in C^{1,1}(\overline{\Omega})$, interior elliptic regularity (Theorem 8.10 in [9]) implies that, for any $\Omega^{\prime}\Subset\Omega\setminus\overline{D}$,
@@
    $\hat{\mathcal{L}}_{2}\mathscr{U}:=\nabla\cdot(\mathscr{A}_{2}\nabla\mathscr{U})+\hat{c}\,\mathscr{U}.$
@@
    $F_{2}(\mathbf{x},t)=\nabla\cdot\big(\mathbf{I}-\mathscr{A}_{2}\big)\nabla u_{0}-c\,u_{0}-\sum_{k=2}^{l_{0}}\alpha_{k}u_{0}^{k}.$ (5.3)
@@
    where the constant $C$ depends only on $d$, $l_{0}$, $D$, $\|\mathscr{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$.
@@
Properties 1 follows immediately from the compact support in $D$ of $\mathbf{I}-\mathscr{A}_{2}(\cdot,t)$, $c(\cdot,t)$, $\alpha_{k}(\cdot,t)$ for $t\in[0,T]$, $k=2,\dots,l_{0}$ combined with the expressions defining $\hat{c}$, $\hat{N}(\mathscr{U})$ and $F_{2}$.
@@
There exists a time $T>0$ ((given below by (5.21))), depending on $\varepsilon$, $d$, $l_{0}$, $D$, $\|\mathscr{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, such that problem (5.1) admits a unique solution $\mathscr{U}$ satisfying:
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{U}+\hat{\mathcal{L}}_{2}\mathscr{U}=F_{2}-\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.6)
@@
where constant $C_{4}(T)$ depends continuously on $T$, $d$, $l_{0}$, $D$, $\|\mathscr{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, but not on $\varepsilon$. Moreover, $C_{4}(T)$ is monotonically increasing in $T$ with $C_{4}(0)<\infty$.
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+\hat{\mathcal{L}}_{2}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+[\partial_{\mathbf{x}}^{\beta},\hat{\mathcal{L}}_{2}]\mathscr{U}=\partial^{\beta}_{\mathrm{x}}F_{2}-\partial^{\beta}_{\mathrm{x}}\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\partial^{\beta}_{\mathrm{x}}\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d},\end{aligned}\right$. (5.8)
where $\partial_{\mathbf{x}}^{\beta},\hat{\mathcal{L}}_{2}]=\partial_{\mathbf{x}}^{\beta}\hat{\mathcal{L}}_{2}-\hat{\mathcal{L}}_{2}\partial_{\mathbf{x}}^{\beta}$ denotes the commutator, which is a differential operator of order at most 2 in spatial derivatives. Multiply the first line in (5.8) with $\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}$, use integration by parts one obtains,
$\mathrm{i}\int_{\mathbb{R}^{d}}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}-\int_{\mathbb{R}^{d}}a^{ij}_{2}(t)\partial_{i}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial_{j}\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ $+\int_{\mathbb{R}^{d}}[\partial_{\mathbf{x}}^{\beta},\hat{\mathcal{L}}_{2}]\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ (5.9)
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{W}+\hat{\mathcal{L}}_{2}\mathscr{W}=-\big(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2})\big)&&\text{in }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{W}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.16)
@@
where $\partial_{\mathbf{x}}^{\gamma},\hat{\mathcal{L}}_{2}]=\partial_{\mathbf{x}}^{\gamma}\hat{\mathcal{L}}_{2}-\hat{\mathcal{L}}_{2}\partial_{\mathbf{x}}^{\gamma}$ denotes the commutator, which is a differential operator of order 0 in spatial derivatives. By using (5.2), the fact that $\mathscr{V}_{1},\mathscr{V}_{2}\in X_{\varepsilon}$ implies $\|\mathscr{V}_{i}\|_{W^{1,\infty}(\mathbb{R}^{d})}\leqslant C\sqrt{\varepsilon}$ for $i=1,2$ and Lemma 5.1, we calculate the term $\partial^{\gamma}_{\mathrm{x}}(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2}))||^{2}_{L^{2}(\mathbb{R}^{d})}$ as follows,